Solve for x: x3 + 149x2 - 50545x - 50693 = 0.

Grade:10

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear ajinkya

x3 + 149x2 - 50545x - 50693 = 0

in cubic equation you always try to find one root by insection.

here in this case x=-1 is its root

so (x+1) is its one factor

now devide the equation by (x+1)

s0

x3 + 149x2 - 50545x - 50693=0

(x+1) ( x2 +148x -50693) =0

now solve 

 x2 +148x -50693 =0

 x = { -148 ±√1482 + 4* 50693} /2

x = 163  and - 311

so final solution is

 x = -1, 163, - 311

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you the answer and detailed
solution very  quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin